OpenStudy (anonymous):

Tracy ran for seven-twelfths hours and Reina ran for one-third hours. How many more hours did Tracy run? one-fourth 1 one-third one-sixth

OpenStudy (anonymous):

I know this! I cant think the answer though!

OpenStudy (anonymous):

7/12 - 4/12 = 3/12 = 1/4 To add/subtract fractions, need to give them the same denominator.

OpenStudy (anonymous):

wha?

OpenStudy (anonymous):

Make 1/3 into 4/12 to be able to subtract from7/12 to find the difference, which is what the question asks.